OpenStudy (anonymous):

A circle is inscribed in a 90-60-30 triangle. How can I find the ratio of the area of the triangle to the circle's area?

OpenStudy (anonymous):

 \[A_t=1/2(x^2\sqrt{3})\] \[A_c=\pi r ^2\]